OpenStudy (anonymous):

calculate no of moles in 1.505*10^23 atoms of ozone?

OpenStudy (anonymous):

@chmvijay

OpenStudy (chmvijay):

sory u dont need molar mass for this 6.023*10^23 atoms makes one mole then 1.505 *10^23 atoms makes = 1*1.505 *10^23 /6.023*10^23

OpenStudy (anonymous):

Ookaaaay...one more Question..calculate the no of atom in 2.5 moles of carbon

OpenStudy (chmvijay):

same thing 1 mole of C has 6.023 *10^23 atoms of C then 2.5 moles of Carbon will have = 2.5 *6.023*10^23